Need help getting y by itself?
2x-3y=9
x+2y-4=0
3x+4y=12
im not sure wether i need to simplfy them or leave them as they are, please help me, thank you

Answer by Alan3354(69443)

Need help getting y by itself?
2x-3y=9
x+2y-4=0
3x+4y=12
im not sure wether i need to simplfy them or leave them as they are, please help me, thank you
----------------------
"getting y by itself" means getting it on one side of the equal sign.
Do what it take to do that.
----
2x-3y=9
Add 3y to both sides.
Divide by 3.
======================
x+2y-4=0
Add 4
subtract x
divide by 2
====================
3x+4y=12
similar to the 2 above.